### v1

# thanks: Corresponding author

Abstract
An accurate determination of singlet–triplet gaps in biradicals, including cyclobutadiene in the automerization barrier region where one has to balance the substantial nondynamical many-electron correlation effects characterizing the singlet ground state with the predominantly dynamical correlations of the lowest-energy triplet, remains a challenge for many quantum chemistry methods. High-level coupled-cluster (CC) approaches, such as the CC method with a full treatment of singly, doubly, and triply excited clusters (CCSDT), are often capable of providing reliable results, but the routine application of such methods is hindered by their high computational costs. We have recently proposed a practical alternative to converging the CCSDT energetics at small fractions of the computational effort, even when electron correlations become stronger and connected triply excited clusters are larger and nonperturbative, by merging the CC($P$;$Q$) moment expansions with the selected configuration interaction methodology abbreviated as CIPSI. We demonstrate that one can accurately approximate the highly accurate CCSDT potential surfaces characterizing the lowest singlet and triplet states of cyclobutadiene along the automerization coordinate and the gap between them using tiny fractions of triply excited cluster amplitudes identified with the help of relatively inexpensive CIPSI Hamiltonian diagonalizations.
I Introduction
Biradicals play a key role in chemistry as reaction intermediates in thermal and photochemical pathwaysMyers et al. (1992); Pedersen et al. (1994); Berson (1978); Zgierski et al. (2005, 2005); Park et al. (2021); Abe et al. (2012) as well as functional materials used in molecular magnets,Dougherty (1991); Malrieu et al. (2014); Cho et al. (2016) battery electrodes,Sun et al. (2014) and organic photovoltaics.Minami and Nakano (2012); Hedley et al. (2017); Niklas and Poluektov (2017); Smith and Michl (2010) An important quantity characterizing the electronic structure of biradicals, especially in the context of designing molecules for magnetic, electrochemical, and photovoltaic applications, is the energy gap $\Delta E_{\text{S--T}}$ separating the lowest-lying singlet and triplet states (throughout this work, we define $\Delta E_{\text{S--T}}$ as $E_{\rm S}-E_{\rm T}$, where $E_{\rm S}$ and $E_{\rm T}$ are the electronic energies of the relevant singlet and triplet states, i.e., when the singlet is lower than the triplet, $\Delta E_{\text{S--T}}<0$). Accurate computational determination of the $\Delta E_{\text{S--T}}$ values in biradicals remains, however, a difficult task because it requires balancing the strong nondynamical many-electron correlation effects characterizing the low-spin singlet states with the predominantly dynamical correlations associated with their high-spin triplet counterparts. Slipchenko and Krylov (2002); Włoch et al. (2007); Li and Paldus (2008); Demel et al. (2008); Saito et al. (2011); Ess et al. (2011); Shen and Piecuch (2012); Abe (2013); Garza et al. (2014); Ibeji and Ghosh (2015); Ajala et al. (2017); Stoneburner et al. (2017); Zimmerman (2017); Shen and Piecuch (2021); Gulania et al. (2021); Boyn and Mazziotti (2021); Chakraborty et al. (2022)
This challenge is exemplified by the cyclobutadiene molecule, which is the focus of the present study and which has fascinated experimental and theoretical chemists for decades with questions surrounding its low-lying electronic states, anti-aromaticity, and reactivity in cycloaddition and isomerization reactions. In the $D_{4\text{h}}$-symmetric square structure corresponding to the barrier along the automerization coordinate or the minimum on the lowest triplet potential, cyclobutadiene is a biradical with its four valence $\pi$ orbitals arranged in a network consisting of the nondegenerate $a_{2u}$ orbital, the doubly degenerate $e_{g}$ shell, and the nondegenerate $b_{1u}$ orbital. The distribution of two of the four valence electrons among the pair of degenerate frontier $e_{g}$ orbitals gives rise to three singlet states of the $B_{1g}(D_{4\text{h}})$, $A_{1g}(D_{4\text{h}})$, and $B_{2g}(D_{4\text{h}})$ symmetries and a $A_{2g}(D_{4\text{h}})$-symmetric triplet state, all involving the doubly occupied $a_{2u}$ and unoccupied $b_{1u}$ orbitals and the partially occupied $e_{g}$ shell in their zeroth-order description. Gulania et al. (2021); Craig (1950); Buenker and Peyerimhoff (1968); Nakamura et al. (1989); Balková and Bartlett (1994); Levchenko and Krylov (2004); Eckert-Maksić et al. (2006); Monino et al. (2022) As shown in the early ab initio calculations, Craig (1950); Buenker and Peyerimhoff (1968); Borden (1975); Kollmar and Staemmler (1977); Borden et al. (1978) and as confirmed in many subsequent theoretical studies, such as those reported in Refs. Gulania et al., 2021; Nakamura et al., 1989; Balková and Bartlett, 1994; Levchenko and Krylov, 2004; Eckert-Maksić et al., 2006; Saito et al., 2011; Ajala et al., 2017; Stoneburner et al., 2017; Zimmerman, 2017; Boyn and Mazziotti, 2021; Monino et al., 2022; Chakraborty et al., 2022, the lowest singlet of the $B_{1g}(D_{4\text{h}})$ symmetry, which has a substantial multiconfigurational character, is the ground state, whereas the predominantly single-reference $A_{2g}(D_{4\text{h}})$-symmetric triplet, in violation of Hund’s rule, is the first excited state. To be more specific, if one orients cyclobutadiene such that the two $C_{2}$ axes bisect the carbon–carbon bonds, which is a convention adopted in the present study, the ${}^{1}B_{1g}(D_{4\text{h}})$ ground state of the square structure is dominated by two closed-shell determinants in which one of the two degenerate $e_{g}$ orbitals is occupied by two electrons and the other one is empty (see, e.g., Refs. Borden et al., 1978; Nakamura et al., 1989; Levchenko and Krylov, 2004; Monino et al., 2022). This should be contrasted with the lowest-energy ${}^{3}A_{2g}(D_{4\text{h}})$ state, which is characterized by single occupancy of each of the $e_{g}$ orbitals.
While the lowest ${}^{3}A_{2g}(D_{\text{4h}})$ state is stable in the square geometry, which is a minimum on the corresponding triplet surface, the ${}^{1}B_{1g}(D_{\text{4h}})$ ground state is unstable with respect to the rectangular distortion of the carbon–carbon bonds that lifts the degeneracy of the valence $e_{g}$ orbitals and lowers its total electronic energy due to the pseudo-Jahn–Teller effect. This results in the formation of the $D_{\text{2h}}$-symmetric rectangular species characterized by the closed-shell, predominantly single-determinantal, ${}^{1}A_{g}(D_{\text{2h}})$ ground state, which represents a minimum on the lowest-energy singlet potential. Gulania et al. (2021); Borden et al. (1978); Nakamura et al. (1989); Balková and Bartlett (1994); Levchenko and Krylov (2004); Boyn and Mazziotti (2021); Monino et al. (2022) The distortion of the multiconfigurational ${}^{1}B_{1g}(D_{4\text{h}})$ state into the ${}^{1}A_{g}(D_{2\text{h}})$ state coincides with the automerization coordinate in cyclobutadiene, which describes the conversion of the rectangular, $D_{2\text{h}}$-symmetric, closed-shell reactant (R) into the equivalent product conformer by passing through the square, $D_{4\text{h}}$-symmetric, biradical transition state (TS). Obtaining an accurate description of the potential energy curves (PECs) characterizing the lowest singlet [${}^{1}A_{g}(D_{2\text{h}})$] and triplet [${}^{3}B_{1g}(D_{2\text{h}})$] states of cyclobutadiene along its $D_{2\text{h}}$-symmetric automerization coordinate, and the gap between them, particularly in the neighborhood of the square biradical species, remains a significant challenge for modern ab initio techniques as it requires a high-level treatment of many-electron correlation effects in order to accurately capture and balance the strong nondynamical correlations associated with the multiconfigurational singlet state with the largely dynamical correlations dominating the triplet state.
A traditional way of addressing this and similar challenges is to use multireference approaches, Roos (1987); Schmidt and Gordon (1998); Szalay et al. (2012); Roca-Sanjuán et al. (2012); Chattopadhyay et al. (2016); Lyakh et al. (2012); Piecuch and Kowalski (2002); Evangelista (2018) but in this work we focus on the single-reference coupled-cluster (CC) methodology,Coester (1958); Coester and Kümmel (1960); Čížek (1966, 1969); Paldus et al. (1972); Paldus and Li (1999); Bartlett and Musiał (2007) which employs the exponential wave function ansatzHubbard (1957); Hugenholtz (1957)
$|\Psi\rangle=e^{T}|\Phi\rangle,$ (1)
where $|\Phi\rangle$ is the $N$-electron reference determinant that serves as a Fermi vacuum and $T=\sum_{n=1}^{N}T_{n}$ is the cluster operator, with $T_{n}$ designating the $n$-body component of $T$ responsible for generating connected $n$-particle–$n$-hole ($n$p-$n$h) excitations out of $|\Phi\rangle$. It is well established that as long as the number of strongly correlated electrons is not too large, the standard hierarchy of CC approximations, including the CC method with singles and doubles (CCSD),Purvis and Bartlett (1982); Cullen and Zerner (1982); Scuseria et al. (1987); Piecuch and Paldus (1989) obtained by truncating $T$ at $T_{2}$, the CC approach with singles, doubles, and triples (CCSDT),Hoffmann and Schaefer (1986); Noga and Bartlett (1987); Scuseria and Schaefer (1988); Watts and Bartlett (1990) in which $T$ is truncated at $T_{3}$, the CC method with singles, doubles, triples, and quadruples (CCSDTQ),Oliphant and Adamowicz (1991); Kucharski and Bartlett (1991, 1992); Piecuch and Adamowicz (1994) where $T$ is truncated at $T_{4}$, and so on, rapidly converges to the exact, full configuration interaction (CI) limit. As a result, the single-reference CC approaches with a full treatment of higher-rank $T_{n}$ clusters with $n>2$, such as CCSDT or CCSDTQ, are often capable of accurately describing multireference situations, including substantial bond rearrangements in the course of chemical reactions and, what is especially important for this study, singlet–triplet gaps in biradical species, by capturing the relevant dynamical and nondynamical correlation effects via particle–hole excitations from a single determinant, without having to involve genuine multireference concepts.
In particular, the calculations reported in Refs. Balková and Bartlett, 1994; Chakraborty et al., 2022; Monino et al., 2022 show that the full treatment of $T_{1}$, $T_{2}$, and $T_{3}$ clusters provided by CCSDT offers a highly accurate description of the total electronic energies of the lowest ${}^{1}B_{1g}(D_{\rm 4h})$ and ${}^{3}A_{2g}(D_{\rm 4h})$ states of the square cyclobutadiene and the gap between them. As demonstrated in Figure 1, this remains true when examining the PECs characterizing the lowest-energy singlet and triplet states of cyclobutadiene along the entire $D_{\rm 2h}$-symmetric automerization reaction path. In determining the lowest-energy ${}^{1}A_{g}(D_{2\text{h}})$ and ${}^{3}B_{1g}(D_{2\text{h}})$ PECs shown in Figure 1 (for information about the electronic structure software used in our calculations, see Section II), we followed the procedure described in Ref. Eckert-Maksić et al., 2006 in which one constructs an approximate, $D_{\rm 2h}$-symmetric, one-dimensional automerization pathway connecting the rectangular minima on the lowest singlet potential via the square TS species by linearly interpolating the carbon–carbon bond distances in cyclobutadiene using the formula
$\ell_{i}(\lambda)=(1-\lambda)\,\ell_{i}(\text{R})+\lambda\,\ell_{i}(\text{TS}),\;\>i=1,2,$ (2)
where $\ell_{1}$ and $\ell_{2}$ are the C–C distances depicted in Figure 1 and $\ell_{i}(\text{R})$ and $\ell_{i}(\text{TS})$ are the carbon–carbon bond lengths characterizing the R and TS structures optimized (along with the C–H distances and H–C–C bond angles) in Ref. Eckert-Maksić et al., 2006 with the multireference average quadratic CC (MR-AQCC) approach.Szalay and Bartlett (1993, 1995) The dimensionless parameter $\lambda$ defining the automerization coordinate varies between 0, corresponding to the R species, and 1, corresponding to the TS structure, going back to 0 (after replacing $\ell_{1}$ by $\ell_{2}$ and vice versa) when the automerization product equivalent to the R species is reached. In the absence of information about the C–H bond lengths and H–C–C bond angles characterizing the intermediate $\lambda=0.2,$ 0.4, 0.6, and 0.8 geometries in Ref. Eckert-Maksić et al., 2006, in determining the lowest ${}^{1}A_{g}(D_{2\text{h}})$ and ${}^{3}B_{1g}(D_{2\text{h}})$ states of these structures, we fixed the C–H distances and H–C–C angles at their values corresponding to the R species.
As shown in Figure 1, the lowest singlet and triplet PECs computed as functions of the $D_{\rm 2h}$-symmetric automerization coordinate $\lambda$ with full CCSDT are in very good agreement with their counterparts obtained with the CI method using perturbative selection made iteratively, Huron et al. (1973); Garniron et al. (2017, 2019) abbreviated as CIPSI, extrapolated to the full CI limit (see Section II for further details). The CCSDT energies are also very close to those determined using the double electron-attachment (DEA) equation-of-motion (EOM) CC methodologyMusiał et al. (2011, 2011); Shen and Piecuch (2013, 2014); Ajala et al. (2017); Gulania et al. (2021); Shen and Piecuch (2021) with a full treatment of 2p and 3p-1h and an active-space treatment of 4p-2h correlations on top of the CCSD description of the underlying closed-shell $({\rm C_{4}H_{4}})^{2+}$ core, Shen and Piecuch (2013, 2014); Ajala et al. (2017); Shen and Piecuch (2021) denoted as DEA-EOMCC(4p-2h)$\{N_{\text{u}}\}$, where $N_{\text{u}}$ designates the number of active unoccupied orbitals of $({\rm C_{4}H_{4}})^{2+}$ included in the calculations to capture the leading 4p-2h effects in the target cyclobutadiene species (to accurately describe the 4p-2h effects associated with the valence orbitals of cyclobutadiene that correlate with the $e_{g}$ and $b_{1u}$ shells of the square TS structure, we set $N_{\text{u}}$ to 3). To illustrate the agreement between full CCSDT, perturbatively corrected and extrapolated CIPSI, and DEA-EOMCC(4p-2h)$\{N_{\text{u}}\}$, which are three independent ab initio methodologies, we compare the vertical $\Delta E_{\text{S--T}}$ values at the R and TS geometries. When using the cc-pVDZDunning (1989) basis set, employed in the calculations reported in Figure 1 and the rest of this article, the CCSDT value of $\Delta E_{\text{S--T}}$ at the TS geometry is $-4.8$ kcal/mol. This is very close to $-5.2$ kcal/mol resulting from the state-of-the-art DEA-EOMCC(4p-2h)$\{N_{\text{u}}\}$ calculations and $-5.9$ kcal/mol obtained with CIPSI. The CCSDT, DEA-EOMCC(4p-2h)$\{N_{\text{u}}\}$, and perturbatively corrected and extrapolated CIPSI values of $\Delta E_{\text{S--T}}$ at the R geometry are $-30.6$, $-30.5$, and $-32.9$ kcal/mol, respectively, again in good agreement with one another.
Given the high accuracy of the lowest singlet and triplet potential surfaces of cyclobutadiene and $\Delta E_{\text{S--T}}$ values along the $D_{\rm 2h}$-symmetric automerization pathway offered by full CCSDT, it may be tempting to turn to the approximate treatments of $T_{3}$ correlations that replace the expensive iterative $\mathscr{N}^{8}$ computational steps of CCSDT, where $\mathscr{N}$ is a measure of the system size, by the more practical $\mathscr{N}^{6}$ operations of CCSD combined with the noniterative $\mathscr{N}^{7}$ steps needed to correct the CCSD energetics for the leading $T_{3}$ correlations, as in the widely used CCSD(T) approach Raghavachari et al. (1989); Watts et al. (1993) or its more robust completely renormalized (CR) CC counterpart abbreviated as CR-CC(2,3).Piecuch and Włoch (2005); Piecuch et al. (2006); Włoch et al. (2006, 2007) Unfortunately, neither CCSD(T) and CR-CC(2,3) nor any of the other noniterative triples corrections to CCSD, such as CCSD(T)Λ,Stanton (1997); Crawford and Stanton (1998); Kucharski and Bartlett (1998) CCSD(2)T,Hirata et al. (2001, 2004); Gwaltney and Head-Gordon (2000, 2001) CR-CCSD(T)Piecuch and Kowalski (2000); Kowalski and Piecuch (2000); Piecuch et al. (2002, 2004) and its locally renormalized extension,Kowalski and Piecuch (2005) $\Lambda$-CCSD(T),Taube and Bartlett (2008, 2008) and CCSD(${\rm T}\mbox{--}n$),Eriksen et al. (2014, 2014) are capable of providing accurate results when the coupling of the lower-order $T_{1}$ and $T_{2}$ components of the cluster operator with their higher-rank $T_{3}$ counterpart becomes large. For example, even the most robust triples correction to CCSD defining CR-CC(2,3), which improves CCSD(T) and other similar approaches in situations involving electronic quasi-degeneracies, such as those present in single bond breaking, Shen and Piecuch (2012); Piecuch and Włoch (2005); Piecuch et al. (2006); Włoch et al. (2006, 2007); Piecuch et al. (2007); Ge et al. (2007, 2008) struggles in describing the PEC of the lowest ${}^{1}A_{g}(D_{2\text{h}})$ state of cyclobutadiene in the neighborhood of the automerization barrier region, where $T_{3}$ clusters become large, nonperturbative, and strongly coupled to $T_{1}$ and $T_{2}$.Shen and Piecuch (2012) As a result, as shown in Ref. Chakraborty et al., 2022 and this study, the CR-CC(2,3) $\Delta E_{\text{S--T}}$ value at the singlet TS structure, of 4.4 kcal/mol when the cc-pVDZ basis is used, is in large error (including incorrect sign) relative to its CCSDT $-4.8$ kcal/mol counterpart [CCSD(T) gives 3.9 kcal/mol, which is similarly inaccurate].
Problems with applying noniterative corrections to CCSD in situations where $T_{n}$ components with $n>2$, such as $T_{3}$, are not only large and nonperturbative, but also strongly coupled to their lower-rank $T_{1}$ and $T_{2}$ counterparts, have motivated us to develop the generalization of the biorthogonal moment expansions, which in the past resulted in the CR-CC approaches, such as CR-CC(2,3) and its excited-state and higher-order extensions, Piecuch and Włoch (2005); Piecuch et al. (2006); Włoch et al. (2006, 2007); Piecuch et al. (2006, 2009); Fradelos et al. (2011); Piecuch et al. (2007, 2008); Horoi et al. (2007); Bauman et al. (2017); Magoulas et al. (2018) to unconventional truncations in the cluster and EOMCCEmrich (1981, 1981); Geertsen et al. (1989); Stanton and Bartlett (1993) excitation operators, designated as CC($P$;$Q$). Shen and Piecuch (2012, 2012, 2012); Bauman et al. (2017); Magoulas et al. (2018); Yuwono et al. (2019); Deustua et al. (2017); Yuwono et al. (2020); Deustua et al. (2021); Chakraborty et al. (2022); Gururangan et al. (2021); Gururangan and Piecuch (2023) By incorporating the dominant contributions to the higher–than–two-body clusters into the iterative steps, so that $T_{1}$ and $T_{2}$ amplitudes can relax compared to their CCSD values when $T_{n}$ components with $n>2$ become more substantial, and correcting the results for the remaining many-electron correlation effects of interest using suitably defined moment expansions, the CC($P$;$Q$) formalism provides us with the opportunity to converge or accurately approximate the parent high-level CCSDT, CCSDTQ, and similar energetics at small fractions of the computational costs, even when noniterative corrections to CCSD fail or struggle. Focusing on full CCSDT, which provides the parent data for the lowest singlet and triplet PECs of cyclobutadiene examined in this work, a few different CC($P$;$Q$) approaches designed to converge CCSDT energetics have been developed. In the initial, active-orbital-based, variant of CC($P$;$Q$), abbreviated as CC(t;3), which is part of the larger CC(t;3), CC(t,q;3), CC(t,q;3,4), CC(q;4), etc. hierarchy,Shen and Piecuch (2012, 2012, 2012); Bauman et al. (2017); Magoulas et al. (2018); Yuwono et al. (2019) the leading $T_{3}$ amplitudes that enter the iterative steps preceding the determination of the CC($P$;$Q$) corrections are obtained using the active-space CCSDt approach.Oliphant and Adamowicz (1992, 1993); Piecuch et al. (1993); Piecuch and Adamowicz (1994, 1995); Ghose et al. (1995); Adamowicz et al. (1998); Piecuch et al. (1999, 1999); Piecuch (2010) In the more black-box semi-stochastic Deustua et al. (2017); Yuwono et al. (2020); Deustua et al. (2021); Chakraborty et al. (2022) and CIPSI-driven CC($P$;$Q$)Gururangan et al. (2021) approaches aimed at converging CCSDT, the dominant triply excited cluster amplitudes are identified with the help of CI Booth et al. (2009); Cleland et al. (2010); Dobrautz et al. (2019); Ghanem et al. (2019, 2020) or CCThom (2010); Franklin et al. (2016); Spencer and Thom (2016); Scott and Thom (2017) Quantum Monte Carlo wave function propagations in the many-electron Hilbert space, in the former case, and the sequences of Hamiltonian diagonalizations constructed in the CIPSI algorithmHuron et al. (1973); Garniron et al. (2017, 2019) in the case of the latter method. In the recently introduced adaptive CC($P$;$Q$) formalism, Gururangan and Piecuch (2023) the triply excited determinants and amplitudes defining the leading $T_{3}$ contributions in the iterative steps of CC($P$;$Q$) calculations are identified using the intrinsic structure of the CC($P$;$Q$) energy corrections.
In this study, we focus on the CIPSI-driven CC($P$;$Q$) methodology of Ref. Gururangan et al., 2021. Our main goal is to answer the question how efficient this methodology is in converging the lowest singlet and triplet potentials of cyclobutadiene along the $D_{\rm 2h}$-symmetric automerization coordinate resulting from the highly accurate CCSDT computations shown in Figure 1. As demonstrated in our initial study,Gururangan et al. (2021) the CIPSI-driven CC($P$;$Q$) approach is capable of producing the near-CCSDT energetics for singlet electronic states using tiny fractions of triply excited cluster amplitudes in the iterative parts of the CC($P$;$Q$) algorithm that are considerably smaller than those used in the analogous semi-stochastic and active-orbital-based CC($P$;$Q$) considerations. One of the key objectives of this work is to determine if similar observations apply to the CIPSI-driven CC($P$;$Q$) calculations for the lowest singlet and triplet potentials of cyclobutadiene along its automerization coordinate and the gap between them. The role of the CC($P$;$Q$) moment corrections in improving the resulting energetics is discussed as well.
II Theory and Computational Details
We begin by summarizing the key ingredients of the CC($P$;$Q$) formalism, as applied to the ground-state problem or, in general, to the lowest state of a given symmetry for which a suitable single-determinantal reference can be found. Each CC($P$;$Q$) calculation requires defining two disjoint subspaces of the many-electron Hilbert space, called the $P$ and $Q$ spaces, designated as $\mathscr{H}^{(P)}$ and $\mathscr{H}^{(Q)}$, respectively. The former space consists of the excited determinants $|\Phi_{K}\rangle=E_{K}|\Phi\rangle$ which, together with the reference function $|\Phi\rangle$, dominate the electronic state of interest ($E_{K}$ is the elementary particle–hole excitation operator that generates $|\Phi_{K}\rangle$ from $|\Phi\rangle$). The determinants spanning the complementary $Q$ space $\mathscr{H}^{(Q)}$ are used to form the noniterative correction $\delta(\mbox{$P$;$Q$})$ which captures higher-order correlation effects the CC calculations in the $P$ space do not describe.
All CC($P$;$Q$) computations consist of two stages. In the first, iterative, stage, denoted as CC($P$), we solve the CC amplitude equations in $\mathscr{H}^{(P)}$ to determine amplitudes $t_{K}$ that define the $P$-space cluster operator
$T^{(P)}=\sum_{|\Phi_{K}\rangle\in\mathscr{H}^{(P)}}t_{K}E_{K}.$ (3)
This is done by employing the conventional projective technique adopted in the majority of single-reference CC calculations, i.e., by solving the system
$\mathfrak{M}_{K}(P)=0,\;|\Phi_{K}\rangle\in\mathscr{H}^{(P)},$ (4)
where
$\mathfrak{M}_{K}(P)=\langle\Phi_{K}|\overline{H}^{(P)}|\Phi\rangle,$ (5)
with $\overline{H}^{(P)}=e^{-T^{(P)}}He^{T^{(P)}}$ representing the similarity-transformed Hamiltonian, are generalized moments of the CC($P$) equations.Piecuch and Kowalski (2000); Kowalski and Piecuch (2000); Jankowski et al. (1991) Once the cluster amplitudes $t_{K}$ defining $T^{(P)}$ are determined, the CC($P$) energy is calculated in a usual way as
$E^{(P)}=\langle\Phi|\overline{H}^{(P)}|\Phi\rangle.$ (6)
In the second stage of the CC($P$;$Q$) procedure, we construct the aforementioned noniterative correction $\delta(\mbox{$P$;$Q$})$ using the expression
$\delta(\mbox{$P$;$Q$})=\sum_{|\Phi_{K}\rangle\in\mathscr{H}^{(Q)}}\ell_{K}(P)\,\mathfrak{M}_{K}(P),$ (7)
where coefficients $\ell_{K}(P)$ multiplying moments $\mathfrak{M}_{K}(P)$ are defined as
$\ell_{K}(P)=\langle\Phi|(1+\Lambda^{(P)})\overline{H}^{(P)}|\Phi_{K}\rangle/D_{K}^{(P)},$ (8)
with $D_{K}^{(P)}=E^{(P)}-\langle\Phi_{K}|\overline{H}^{(P)}|\Phi_{K}\rangle$ designating the relevant Epstein–Nesbet-like denominators. The hole–particle deexcitation operator
$\Lambda^{(P)}=\sum_{|\Phi_{K}\rangle\in\mathscr{H}^{(P)}}\lambda_{K}(E_{K})^{\dagger}$ (9)
in Eq. (8), which defines the bra state $\langle\tilde{\Psi}^{(P)}|=\langle\Phi|(1+\Lambda^{(P)})e^{-T^{(P)}}$ matching the CC($P$) ket state $|\Psi^{(P)}\rangle=e^{T^{(P)}}|\Phi\rangle$, is obtained by solving the linear system
$\langle\Phi|(1+\Lambda^{(P)})\overline{H}^{(P)}|\Phi_{K}\rangle=E^{(P)}\lambda_{K},\;|\Phi_{K}\rangle\in\mathscr{H}^{(P)}.$ (10)
The final CC($P$;$Q$) energy is obtained using the formula
$E^{(P+Q)}=E^{(P)}+\delta(\mbox{$P$;$Q$}).$ (11)
One of the main advantages of the CC($P$;$Q$) methodology is its flexibility. In particular, we can make a wide variety of conventional as well as unconventional choices of the $P$ and $Q$ spaces, adjusting them to the nature of the electronic states of interest and adopting different numerical procedures in their construction. Conventional choices for the $P$ and $Q$ spaces, based on the many-body ranks of the determinants included in them, result in the left-eigenstate CR-CC methods, such as the CR-CC(2,3) approach discussed in the Introduction, in which the former space consists of all singly and doubly excited determinants and the latter space is spanned by all triples. We can, however, also make unconventional choices, including those adopted in the CC(t;3), CC(t,q;3), CC(t,q;3,4), CC(q;4), etc. hierarchy Shen and Piecuch (2012, 2012, 2012); Bauman et al. (2017); Magoulas et al. (2018); Yuwono et al. (2019) and the semi-stochastic, Deustua et al. (2017); Yuwono et al. (2020); Deustua et al. (2021); Chakraborty et al. (2022) adaptive,Gururangan and Piecuch (2023) and CIPSI-drivenGururangan et al. (2021) CC($P$;$Q$) methods, mentioned in the Introduction as well, in which the suitably chosen subsets of higher–than–doubly excited determinants are incorporated into the underlying $P$ spaces, in addition to all singles and doubles, to relax the lower-rank $T_{1}$ and $T_{2}$ clusters in the presence of their higher-rank counterparts, such as the leading $T_{3}$ contributions, which the CCSD(T), CR-CC(2,3), $\Lambda$-CCSD(T), and similar approaches are not designed to do. Having some higher–than–doubly excited determinants in the $P$ space provides us with a straightforward and computationally efficient mechanism to account for the coupling between the lower- and higher-order components of the cluster operator, which cannot be neglected when $T_{n}$ contributions with $n>2$, such as $T_{3}$, become large and nonperturbative, as is the case when the automerization barrier region of the lowest-energy singlet potential of cyclobutadiene is examined. This, in turn, allows us to recover the full CCSDT, CCSDTQ, and similar energetics without running into the very expensive, often prohibitive, computational costs associated with the high-level CC methods of this type.
In the case of the CIPSI-driven CC($P$;$Q$) approach, introduced in Ref. Gururangan et al., 2021 and investigated in this study, the desired subsets of higher–than–doubly excited determinants incorporated into the underlying $P$ spaces are identified with the help of sequences of relatively inexpensive Hamiltonian diagonalizations in systematically grown, recursively defined, subspaces of the many-electron Hilbert space, denoted as $\mathscr{V}_{\text{int}}^{(k)}$, where $k=0,1,2,\ldots$ enumerates the consecutive CIPSI iterations. In doing so, we follow the CIPSI algorithm, originally proposed in Ref. Huron et al., 1973, further developed in Refs. Garniron et al., 2017, 2019, and available in the Quantum Package 2.0 software.Garniron et al. (2019) Given our interest in using CIPSI, which is one of the selected CI approachesWhitten and Hackmeyer (1969); Bender and Davidson (1969); Huron et al. (1973); Buenker and Peyerimhoff (1974) (see Refs. Schriber and Evangelista, 2016, 2017; Tubman et al., 2016, 2020; Liu and Hoffmann, 2016; Zhang et al., 2020; Holmes et al., 2016; Sharma et al., 2017; Li et al., 2018 for other examples), within the single-reference CC($P$;$Q$) framework, the initial subspaces $\mathscr{V}_{\text{int}}^{(0)}$ adopted in our work are always spanned by the restricted Hartree–Fock (RHF) or restricted open-shell Hartree–Fock (ROHF) determinants. Once $\mathscr{V}_{\text{int}}^{(0)}$ is defined, each subsequent subspace $\mathscr{V}_{\text{int}}^{(k+1)}$ with $k\geq 0$ is constructed by enlarging its $\mathscr{V}_{\text{int}}^{(k)}$ predecessor with the subset of the leading singly and doubly exited determinants generated out of it, identified with the help of the many-body perturbation theory (MBPT). Thus, if $|\Psi_{k}^{(\text{CIPSI})}\rangle=\sum_{|\Phi_{I}\rangle\in\mathscr{V}_{\text{int}}^{(k)}}c_{I}|\Phi_{I}\rangle$ and $E_{\text{var},k}$ are the CI wave function and energy obtained in $\mathscr{V}_{\text{int}}^{(k)}$, and if the space of all singles and doubles out of $|\Psi_{k}^{(\text{CIPSI})}\rangle$ is designated as $\mathscr{V}_{\text{ext}}^{(k)}$, the subset of determinants $|\Phi_{\alpha}\rangle\in\mathscr{V}_{\text{ext}}^{(k)}$ selected for inclusion in $\mathscr{V}_{\text{int}}^{(k+1)}$ consists of those that have the largest $e_{\alpha,k}^{(2)}=|\langle\Phi_{\alpha}|H|\Psi_{k}^{\text{(CIPSI)}}\rangle|^{2}/(E_{\text{var},k}-\langle\Phi_{\alpha}|H|\Phi_{\alpha}\rangle)$ contributions to the perturbative correction $\Delta E_{k}^{(2)}=\sum_{|\Phi_{\alpha}\rangle\in\mathscr{V}_{\text{ext}}^{(k)}}e_{\alpha,k}^{(2)}$ to $E_{\text{var},k}$. Their selection is accomplished by arranging the sampled determinants $|\Phi_{\alpha}\rangle\in\mathscr{V}_{\text{ext}}^{(k)}$ in descending order according to their $|e_{\alpha,k}^{(2)}|$ values and enlarging $\mathscr{V}_{\text{int}}^{(k)}$, determinant by determinant, starting with the $|\Phi_{\alpha}\rangle$s associated with the largest $|e_{\alpha,k}^{(2)}|$ contributions and moving toward those characterized by smaller values of $|e_{\alpha,k}^{(2)}|$, until the dimension of $\mathscr{V}_{\text{int}}^{(k+1)}$ exceeds that of its $\mathscr{V}_{\text{int}}^{(k)}$ predecessor by a user-defined factor $f>1$, which in all the calculations performed in this study was set to its default value of 2 [the actual number of determinants included in $\mathscr{V}_{\text{int}}^{(k+1)}$ is usually slightly larger than $f$ times the dimension of $\mathscr{V}_{\text{int}}^{(k)}$ since one may have to add extra determinants in $\mathscr{V}_{\text{int}}^{(k+1)}$ to make sure that the corresponding CI wave function $|\Psi_{k+1}^{(\text{CIPSI})}\rangle$ is an eigenstate of the total spin $S^{2}$ and $S_{z}$ operators]. To reduce the computational costs associated with the above procedure of enlarging the $\mathscr{V}_{\text{int}}^{(k)}$ space to obtain $\mathscr{V}_{\text{int}}^{(k+1)}$, in all the CIPSI-driven CC($P$;$Q$) computations reported in this work, we relied on a semi-stochastic version of the above determinant selection algorithm implemented in Quantum Package 2.0, in which one stochastically filters out the most important singly and doubly excited determinants out of $|\Psi_{k}^{(\text{CIPSI})}\rangle$, so that only a small subset of singles and doubles ends up in the $\mathscr{V}_{\text{ext}}^{(k)}$ space prior to determining and analyzing the $e_{\alpha,k}^{(2)}$ contributions. The $e_{\alpha,k}^{(2)}$ values, in addition to guiding the process of enlarging diagonalization spaces $\mathscr{V}_{\text{int}}^{(k)}$ and allowing us to evaluate the perturbatively corrected CIPSI energies $E_{\text{var},k}+\Delta E_{k}^{(2)}$, can be used to calculate the renormalized second-order corrections $\Delta E_{\text{r},k}^{(2)}$ introduced in Ref. Garniron et al., 2019 and the $E_{\text{var},k}+\Delta E_{\text{r},k}^{(2)}$ energies.
To produce the final wave function $|\Psi^{(\text{CIPSI})}\rangle$, needed to construct the list of higher–than–doubly excited determinants to be included in the $P$ space of a given CIPSI-driven CC($P$;$Q$) calculation, and determine the associated variational ($E_{\text{var}}$) and perturbatively corrected [$E_{\text{var}}+\Delta E^{(2)}$ or $E_{\text{var}}+\Delta E_{\text{r}}^{(2)}$] CIPSI energies, the sequence of Hamiltonian diagonalizations defining the underlying CIPSI run must be terminated. This could be done by stopping at the first iteration $k$ for which the absolute value of the second-order MBPT correction $\Delta E_{k}^{(2)}$ falls below a user-defined threshold $\eta$, but, given our interest in examining the convergence of the CIPSI-based CC($P$;$Q$) calculations toward the desired high-level CC energetics, represented in this study by CCSDT, using systematically grown $P$ spaces obtained with the help of CIPSI, in this work we follow Ref. Gururangan et al., 2021 and stop when the number of determinants in the diagonalization space equalizes or exceeds the user-defined parameter $N_{\text{det(in)}}$. To ensure that the CIPSI sequences preceding our CC($P$;$Q$) calculations did not terminate too soon, before the dimensions of terminal diagonalization spaces became greater than or equal to $N_{\text{det(in)}}$, we set the aforementioned parameter $\eta$ to 1 microhartree. As a result [putting aside the convergence threshold used in the CC($P$) iterations, which we set to $10^{-8}$ hartree], all CIPSI-driven CC($P$;$Q$) computations reported in this article, along with the underlying $P$ spaces, were controlled by a single input variable $N_{\text{det(in)}}$. In addition to $N_{\text{det(in)}}$, in presenting our CC($P$) and CC($P$;$Q$) results for the lowest singlet and triplet potentials of cyclobutadiene, we also provide information about the numbers of determinants included in the terminal CIPSI wave functions $|\Psi^{(\text{CIPSI})}\rangle$ obtained for various values of $N_{\text{det(in)}}$, designated as $N_{\text{det(out)}}$. Given our choice of the subspace enlargement parameter $f$, the $N_{\text{det(out)}}$ values characterizing the $|\Psi^{(\text{CIPSI})}\rangle$ states used to identify the triply excited determinants for inclusion in the $P$ spaces employed in our CIPSI-driven CC($P$;$Q$) computations were always between $N_{\text{det(in)}}$ and $2N_{\text{det(in)}}$. With all of this in mind, the algorithm used in the CIPSI-enabled CC($P$;$Q$) calculations reported in this study, aimed at converging the CCSDT energetics, can be summarized as follows: Gururangan et al. (2021)
  1. 1.
    Choose a wave function termination parameter $N_{\text{det(in)}}$ and execute a CIPSI diagonalization sequence starting from the one-dimensional subspace $\mathscr{V}_{\text{int}}^{(0)}$ spanned by the RHF or ROHF reference determinant $|\Phi\rangle$ to obtain the $|\Psi^{\text{(CIPSI)}}\rangle$ state.
  2. 2.
    Extract the list of triply excited determinants included in $|\Psi^{\text{(CIPSI)}}\rangle$ and combine it with all singly and doubly excited determinants relative to $|\Phi\rangle$ to obtain the $P$ space for CC($P$;$Q$) calculations.
  3. 3.
    Solve the CC($P$) amplitude equations, Eq. (4), to determine the cluster operator $T^{(P)}=T_{1}+T_{2}+T_{3}^{\text{(CIPSI)}}$, where $T_{3}^{\text{(CIPSI)}}$ is the three-body component of $T^{(P)}$ defined using the list of triply excited determinants extracted from $|\Psi^{\text{(CIPSI)}}\rangle$, and energy $E^{(P)}$, Eq. (6). Solve the left-eigenstate CC($P$) system given by Eq. (10) to obtain the companion hole–particle deexcitation operator $\Lambda^{(P)}=\Lambda_{1}+\Lambda_{2}+\Lambda_{3}^{\text{(CIPSI)}}$ in which the triples entering $\Lambda_{3}^{\text{(CIPSI)}}$ are the same as those included in $T_{3}^{\text{(CIPSI)}}$.
  4. 4.
    Calculate the noniterative correction $\delta(\mbox{$P$;$Q$})$, Eq. (7), in which the $Q$ space is defined as the remaining triply excited determinants absent in $|\Psi^{\text{(CIPSI)}}\rangle$, and add it to $E^{(P)}$ to obtain the CC($P$;$Q$) energy $E^{(P+Q)}$, Eq. (11).
The above steps 1–4 can be repeated by increasing $N_{\text{det(in)}}$. The entire process can be stopped when the difference between consecutive $E^{(P+Q)}$ values falls below some small, user-specified, convergence threshold.
In order to perform the CIPSI-driven CC($P$) and CC($P$;$Q$) calculations for the singlet and triplet PECs of cyclobutadiene along the automerization coordinate investigated in this work and examine their convergence toward the CCSDT potentials shown in Figure 1, we used the computer programs described in Ref. Gururangan et al., 2021, interfaced with the RHF, ROHF, and integral transformation routines in GAMESS,Schmidt et al. (1993); Barca et al. (2020); Zahariev et al. (2023) which take advantage of our highly efficient, automatically generated, Fortran CC codes that were previously exploited in implementing the active-orbital-basedShen and Piecuch (2012, 2012, 2012); Bauman et al. (2017); Magoulas et al. (2018) and semi-stochasticDeustua et al. (2017); Yuwono et al. (2020); Deustua et al. (2021) CC($P$;$Q$) approaches. We also used our group’s computer-generated CC codes interfaced with GAMESS to produce the parent CCSDT data. The results of the DEA-EOMCC(4p-2h)$\{N_{\text{u}}\}$ calculations discussed in the Introduction, reported in Figure 1 as well, were carried out with the highly efficient DEA-EOMCC routines developed in Ref. Shen and Piecuch, 2021, which became part of the official GAMESS distribution last year.
As already alluded to above, the lists of triply excited determinants that were needed to construct the $P$ spaces for the CC($P$) and CC($P$;$Q$) calculations corresponding to the various values of the input parameter $N_{\text{det(in)}}$ were extracted from the terminal CIPSI wave functions $|\Psi^{(\text{CIPSI})}\rangle$ obtained with Quantum Package 2.0, whereas the complementary $Q$ spaces, used to determine the $\delta(\mbox{$P$;$Q$})$ corrections, consisted of the remaining triples absent in the $|\Psi^{(\text{CIPSI})}\rangle$ states. We also used Quantum Package 2.0 to obtain the $E_{\text{var}}$, $E_{\text{var}}+\Delta E^{(2)}$, and $E_{\text{var}}+\Delta E_{\text{r}}^{(2)}$ energies associated with the CIPSI runs that provided the lists of triples for the CC($P$) iterations. Our CIPSI-driven CC($P$) and CC($P$;$Q$) methods targeting CCSDT that rely on the triples lists extracted from the CIPSI calculations performed with Quantum Package 2.0 have recently been incorporated in our open-source CCpy package available on GitHub,CCp implemented using a hybrid Python-Fortran programming approach.
To obtain the desired insights into the convergence of the singlet and triplet potentials of cyclobutadiene toward their CCSDT counterparts, we adopted the strategy used in Ref. Gururangan et al., 2021. Thus, for each nuclear geometry along the automerization pathway considered in this work, we carried out a series of CIPSI-driven CC($P$) and CC($P$;$Q$) calculations using the $P$ and $Q$ spaces derived from the increasingly large CIPSI wave functions obtained by varying $N_{\text{det(in)}}$ in an approximately semi-logarithmic manner. We started with $N_{\text{det(in)}}=1$, where the CIPSI wave functions $|\Psi^{(\text{CIPSI})}\rangle$ are the single determinants defining the reference states $|\Phi\rangle$ used in our CC computations (RHF in the case of the singlet and ROHF in the triplet case) and the resulting CC($P$) and CC($P$;$Q$) energies become identical to those obtained with CCSD and CR-CC(2,3), respectively, and went all the way to $N_{\text{det(in)}}=10,000,000$, to reflect on the fact that as the input variable $N_{\text{det(in)}}$ becomes increasingly large and the CIPSI wave functions capture more and more triply excited determinants, the CC($P$;$Q$) energies $E^{(P+Q)}$ approach their CCSDT parents (becoming identical to them when all triply excited determinants are captured by the $|\Psi^{(\text{CIPSI})}\rangle$ states). The energies obtained in the CC($P$) computations approach their CCSDT counterparts as well, but, as discussed in the next section, their convergence toward CCSDT is much slower than that observed in the CC($P$;$Q$) runs. The ability of the CIPSI-driven CC($P$;$Q$) methodology to generate the CCSDT-level energetics using tiny fractions of triply excited determinants in the underlying $P$ spaces captured by the relatively small Hamiltonian diagonalizations, observed in the calculations for the lowest ${}^{1}A_{g}(D_{2\text{h}})$ and ${}^{3}B_{1g}(D_{2\text{h}})$ potentials of cyclobutadiene reported in this study, results in enormous savings in the computational effort compared to CCSDT. Indeed, as explained in Ref. Gururangan et al., 2021, the CIPSI runs using smaller $N_{\text{det(in)}}$ values are much faster than those needed to reach convergence, the CC($P$) calculations using tiny fractions of triples in the $P$ space are orders of magnitude faster than the corresponding CCSDT computations (cf. Ref. Gururangan and Piecuch, 2023 for the examples of timings), and the effort involved in obtaining the noniterative $\delta(\mbox{$P$;$Q$})$ corrections is similar to the determination of the triples corrections of CR-CC(2,3) or CCSD(T).
While our numerical evidence clearly demonstrates that the CIPSI computations characterized by $N_{\text{det(in)}}=10,000,000$ result in the $P$ spaces which are unnecessarily large for accurately approximating the CCSDT energetics using the CIPSI-driven CC($P$;$Q$) approach, we include them in our analysis since they also allowed us to extrapolate the near-full-CI $E_{\text{var}}+\Delta E_{r}^{(2)}$ potentials shown in Figure 1 by following the procedure described in Refs. Garniron et al., 2019; Loos et al., 2020; Monino et al., 2022 (see, also, Ref. Gururangan et al., 2021). In this procedure, the $E_{\text{var},k}+\Delta E_{r,k}^{(2)}$ energies extracted from the last four to six Hamiltonian diagonalizations of the CIPSI sequence leading to the final $|\Psi^{(\text{CIPSI})}\rangle$ state are plotted against the corresponding $\Delta E_{r,k}^{(2)}$ corrections and the resulting data, fit to a line, are extrapolated to the $\Delta E_{r}^{(2)}=0$ limit. In the case of the CIPSI calculations performed in this work, to extrapolate the reasonably smooth PECs for the lowest singlet and triplet states of cyclobutadiene out of our largest CIPSI runs corresponding to $N_{\text{det(in)}}=10,000,000$, shown in Figure 1, we had to use the last six $E_{\text{var},k}+\Delta E_{r,k}^{(2)}$ and $\Delta E_{r,k}^{(2)}$ values of each of these runs. Using fewer than six values resulted in unphysical bumps in the extrapolated potentials. In addition to allowing us to comment on the quality of the CCSDT energetics in the Introduction, the extrapolated CIPSI PECs shown in Figure 1 serve in this study as reference potentials for assessing the accuracy of the $E_{\text{var}}$, $E_{\text{var}}+\Delta E^{(2)}$, and $E_{\text{var}}+\Delta E_{r}^{(2)}$ energies obtained in the CIPSI calculations using various values of $N_{\text{det(in)}}$ from a 1–10,000,000 range.
In generating the results of the CIPSI-driven CC($P$;$Q$) computations for the lowest singlet and triplet states of cyclobutadiene along its automerization coordinate at different values of the wave function terminating parameter $N_{\text{det(in)}}$, discussed in the next section, we adopted the same philosophy as that used in the case of the DEA-EOMCC(4p-2h)$\{N_{\text{u}}\}$, extrapolated $E_{\text{var}}+\Delta E_{r}^{(2)}$, and parent CCSDT potentials shown in Figure 1. Thus, we used Eq. (2), in which the geometries of the R and TS structures on the singlet potential obtained with MR-AQCC were taken from Ref. Eckert-Maksić et al., 2006, to set up a one-dimensional, $D_{\rm 2h}$-symmetric, automerization pathway parameterized by dimensionless variable $\lambda\in[0,1]$. In analogy to the PECs shown in Figure 1, we carried out our CIPSI-based CC($P$;$Q$) and the underlying CC($P$) calculations for the lowest ${}^{1}A_{g}(D_{2\text{h}})$ and ${}^{3}B_{1g}(D_{2\text{h}})$ states of cyclobutadiene and determined the corresponding $E_{\text{var}}$, $E_{\text{var}}+\Delta E^{(2)}$, and $E_{\text{var}}+\Delta E_{\text{r}}^{(2)}$ CIPSI energies for $\lambda=0$, 0.2, 0.4, 0.6, 0.8, and 1, reflecting the resulting PEC segments about $\lambda=1$ to obtain the potentials that connect the rectangular reactant and product minima via the square TS. As in the case of the full CCSDT, perturbatively corrected and extrapolated CIPSI, and DEA-EOMCC(4p-2h)$\{N_{\text{u}}\}$ calculations reported in Figure 1, all of our CIPSI-driven CC($P$) and CC($P$;$Q$) computations were performed using the cc-pVDZ basis set and the four core molecular orbitals correlating with the 1s shells of the carbon atoms were frozen in the post-RHF/ROHF steps. Consistent with the overall symmetry of the automerization pathway examined in this work, the $D_{\rm 2h}$ point group was adopted throughout. In particular, the $P$ and $Q$ spaces used in our CC($P$;$Q$) calculations for the lowest-energy singlet PEC consisted of the $S_{z}=0$ determinants of the $A_{g}(D_{\rm 2h})$ symmetry. In the case of the lowest-energy triplet potential, they consisted of the $S_{z}=1$ $B_{1g}(D_{\rm 2h})$ determinants.
III Results and Discussion
As explained in the Introduction, the primary objective of this study is to examine efficiency of the CIPSI-driven CC($P$;$Q$) methodology in converging the full CCSDT data for the lowest singlet and triplet potentials of cyclobutadiene along its automerization coordinate and the gap between them. We are especially interested in investigating how effective the CIPSI-driven CC($P$;$Q$) approach is in balancing the substantial nondynamical many-electron correlation effects characterizing the lowest ${}^{1}A_{g}(D_{\rm 2h})$ state in the neighborhood of the automerization barrier region, where $T_{3}$ clusters become large, nonperturbative, and strongly coupled to their lower-rank $T_{1}$ and $T_{2}$ counterparts, with the predominantly dynamical correlations characterizing the lowest ${}^{3}B_{1g}(D_{\rm 2h})$ state. As pointed out in the Introduction, using comparisons with the DEA-EOMCC(4p-2h)$\{N_{\text{u}}\}$ and perturbatively corrected and extrapolated CIPSI results, the CCSDT approach, despite its intrinsically single-reference character, captures essentially all relevant many-electron correlation effects needed to accurately describe the lowest singlet and triplet potentials of cyclobutadiene and the separation between them. The question is if the CIPSI-driven CC($P$;$Q$) computations using compact wave functions $|\Psi^{(\text{CIPSI})}\rangle$, resulting from the relatively inexpensive CIPSI diagonalization sequences characterized by the $N_{\text{det(out)}}$ values that are much smaller than the numbers of all $T_{3}$ amplitudes, and tiny fractions of the triply excited determinants in the underlying $P$ spaces are capable of accomplishing the same. We also examine how effective the noniterative $\delta(\mbox{$P$;$Q$})$ corrections are in accelerating convergence of the CC($P$) energetics toward their CCSDT parents and how the rate of convergence of the CIPSI-driven CC($P$;$Q$) calculations toward CCSDT with $N_{\text{det(in)}}$ compares to the analogous convergence of the perturbatively corrected CIPSI energies toward their extrapolated $E_{\text{var}}+\Delta E_{\text{r}}^{(2)}$ values. The results of our CIPSI-driven CC($P$) and CC($P$;$Q$) calculations for the lowest ${}^{1}A_{g}(D_{\rm 2h})$ and ${}^{3}B_{1g}(D_{\rm 2h})$ potentials and $\Delta E_{\text{S--T}}$ gaps of cyclobutadiene, along with the associated $E_{\text{var}}$, $E_{\text{var}}+\Delta E^{(2)}$, and $E_{\text{var}}+\Delta E^{(2)}_{r}$ data, can be found in Tables 13. The CC($P$) and CC($P$;$Q$) potentials and $\Delta E_{\text{S--T}}$ values are also shown in Figure 2.
We begin by commenting on the calculations using $N_{\text{det(in)}}=1$, which help us appreciate the need for including the leading triply excited determinants in the $P$ spaces employed in the CC($P$) and CC($P$;$Q$) computations, especially when $T_{3}$ effects and the coupling between $T_{1}$ and $T_{2}$ clusters and their higher-rank $T_{3}$ counterpart become significant. As explained in Section II, when $N_{\text{det(in)}}=1$, the CIPSI-driven CC($P$) and CC($P$;$Q$) approaches become equivalent to CCSD and CR-CC(2,3), respectively, i.e., one solves the CCSD equations for $T_{1}$ and $T_{2}$ clusters, as if they were decoupled from their higher-rank $T_{3}$ counterpart, and corrects the resulting CCSD energies for the effects of connected triples using the CR-CC(2,3) method. Upon examining the $N_{\text{det(in)}}=1$ CC($P$) values in Table 1, we observe that the $T_{3}$ correlation effects characterizing the lowest-energy ${}^{1}A_{g}(D_{2\text{h}})$ potential, estimated by forming differences between the CCSDT and CCSD energies, are not only large, but also dramatically changing along the automerization coordinate, from $-26.827$ millihartree for the $\lambda=0$ R species to $-47.979$ millihartree when the square TS structure corresponding to $\lambda=1$ is considered. As indicated by the $N_{\text{det(in)}}=1$ CC($P$;$Q$) results shown in Table 1, the incorporation of $T_{3}$ correlations using the noniterative CR-CC(2,3) corrections to the CCSD energies helps, reducing the 26.827, 27.964, 29.667, 32.473, 37.662, and 47.979 millihartree errors relative to CCSDT obtained at $\lambda=0$, 0.2, 0.4, 0.6, 0.8, and 1 with CCSD to 0.848, 1.253, 2.021, 3.582, 7.008, and 14.636 millihartree, respectively, but substantial discrepancies between the CR-CC(2,3) and CCSDT data, especially in the neighborhood of the barrrier region, where they are as large as 7–15 millihartree when $\lambda\in[0.8,1]$, remain. As a result, the quality of the ${}^{1}A_{g}(D_{2\text{h}})$ potential obtained in the CR-CC(2,3) calculations, which is characterized by a large, 13.788 millihartree, nonparallelity error (NPE) relative to its CCSDT parent and which is shown in Figure 2 (b), is poor. Failure of CR-CC(2,3) and, as demonstrated, for example, in Refs. Shen and Piecuch, 2012; Lyakh et al., 2011, of other noniterative triples corrections to CCSD, including CCSD(T)Shen and Piecuch (2012); Lyakh et al. (2011) and CCSD(2)T,Shen and Piecuch (2012) to accurately describe the lowest ${}^{1}A_{g}(D_{2\text{h}})$ state of cyclobutadiene in the automerization barrier region is, in significant part, a consequence of the inability of all such methods to capture the coupling of the lower-rank $T_{1}$ and $T_{2}$ clusters with $T_{3}$, which in the vicinity of the TS geometry, where $T_{3}$ effects are large and nonperturbative, is big enough to substantially alter $T_{1}$ and $T_{2}$ amplitudes compared to their CCSD values. This means that to improve the quality of the ${}^{1}A_{g}(D_{2\text{h}})$ potential obtained with CR-CC(2,3), one should relax $T_{1}$ and $T_{2}$ clusters, adjusting them to the presence of $T_{3}$ correlations, prior to determining noniterative triples corrections. The CIPSI-driven CC($P$;$Q$) methodology allows us to do it in a computationally efficient manner, avoiding expensive CCSDT iterations, by incorporating the leading triply excited determinants identified with the help of the CIPSI runs using sufficiently large $N_{\text{det(in)}}>1$ values into the underlying $P$ spaces and correcting the resulting CC($P$) energies for the remaining $T_{3}$ effects using the $\delta(\mbox{$P$;$Q$})$ corrections.
The situation with the lowest-energy ${}^{3}B_{1g}(D_{\rm 2h})$ potential is different. In this case, as shown in Table 2, the $T_{3}$ correlation effects, estimated, after subtracting the CCSD energies from their CCSDT counterparts, at about $(-25)-(-24)$ millihartree, barely vary with the automerization coordinate $\lambda$ and are accurately described by the CR-CC(2,3) approach, which reproduces the parent CCSDT energetics to within 60 microhartree across the entire ${}^{3}B_{1g}(D_{\rm 2h})$ PEC. Because of this very different behavior of CR-CC(2,3) compared to the lowest-energy singlet potential, which can be seen by comparing the $N_{\text{det(in)}}=1$ CC($P$;$Q$) PECs in the (b) and (d) panels of Figure 2, one ends up with a highly unbalanced description of the lowest singlet and triplet states of cyclobutadiene by the CR-CC(2,3) method, especially in vicinity of the barrier on the ${}^{1}A_{g}(D_{2\text{h}})$ potential. This is reflected in the errors characterizing the singlet–triplet gap values obtained with CR-CC(2,3), relative to CCSDT, which grow from $0.553$ kcal/mol at $\lambda=0$ to $9.222$ kcal/mol when the $\lambda=1$ square TS structure is considered [see the $N_{\text{det(in)}}=1$ CC($P$;$Q$) values of $\Delta E_{\text{S--T}}$ in Table 3; cf., also, Figure 2 (f)]. Once again, the main problem resides in the neglect of the coupling between $T_{1}$ and $T_{2}$ clusters and their higher-rank $T_{3}$ counterpart in the CR-CC(2,3) approach, which results in a poor description of the ${}^{1}A_{g}(D_{2\text{h}})$ potential in the vicinity of the TS geometry that propagates into the similarly poor $\Delta E_{\text{S--T}}$ gap values. To bring the results closer to those obtained with CCSDT, the input variable $N_{\text{det(in)}}$ that controls the CIPSI runs preceding the CC($P$) and CC($P$;$Q$) steps must be increased. The results of the CIPSI-driven CC($P$) and CC($P$;$Q$) computations for the lowest ${}^{1}A_{g}(D_{2\text{h}})$ and ${}^{3}B_{1g}(D_{\rm 2h})$ potentials of cyclobutadiene and the gap between them using representative $N_{\text{det(in)}}>1$ values are discussed next.
As shown in Table 1 and Figure 2 (b), the CC($P$;$Q$) computations for the lowest ${}^{1}A_{g}(D_{2\text{h}})$ state, using CIPSI Hamiltonian diagonalizations to identify the leading triply excited determinants for inclusion in the underlying $P$ spaces, display fast convergence toward CCSDT with $N_{\text{det(in)}}$, independent of the value of $\lambda$. With as little as 101,361101361101,361101 , 361196,965196965196,965196 , 965 $S_{z}=0$ determinants of the $A_{g}(D_{\rm 2h})$ symmetry in the terminal $|\Psi^{(\text{CIPSI})}\rangle$ wave functions generated by the inexpensive CIPSI runs using $N_{\text{det(in)}}=100,000$, which capture tiny fractions, on the order of 0.1–0.2%, of the $14,483,876$ $A_{g}(D_{\rm 2h})$-symmetric $S_{z}=0$ triples, the CC($P$;$Q$) method reduces the 0.848, 1.253, 2.021, 3.582, 7.008, and 14.636 millihartree errors relative to CCSDT obtained at $\lambda=0$, 0.2, 0.4, 0.6, 0.8, and 1 with CR-CC(2,3) to 0.431, 0.552, 0.776, 1.235, 0.628, and 3.539 millihartree, respectively. With the relatively small additional effort corresponding to $N_{\text{det(in)}}=250,000$, which results in 394,080394080394,080394 , 080449,753449753449,753449 , 753 $S_{z}=0$ determinants of the $A_{g}(D_{\rm 2h})$ symmetry in the final CIPSI diagonalization spaces and only 0.5–0.6% of all triples in the underlying $P$ spaces, the differences between the CC($P$;$Q$) and CCSDT energies of the lowest singlet state of cyclobutadiene at $\lambda=0$, 0.2, 0.4, 0.6, 0.8, and 1 decrease to 0.278, 0.272, 0.271, 0.300, 0.336, and 0.458 millihartree, respectively. Clearly, these are massive error reductions compared to the CR-CC(2,3) computations, especially in the barrier region, which highlight the effectiveness of our CIPSI-driven CC($P$;$Q$) strategy and the importance of relaxing $T_{1}$ and $T_{2}$ amplitudes in the presence of the leading $T_{3}$ contributions compared to their CCSD values prior to determining the noniterative corrections for the remaining $T_{3}$ effects. Similar comments apply to the improvements in the troublesome 13.788 millihartree NPE relative to CCSDT characterizing the ${}^{1}A_{g}(D_{2\text{h}})$ potential obtained in the CR-CC(2,3) calculations offered by the CIPSI-driven CC($P$;$Q$) runs. When the CC($P$;$Q$) approach using $N_{\text{det(in)}}=100,000$ is employed, the NPE relative to CCSDT characterizing the resulting ${}^{1}A_{g}(D_{2\text{h}})$ potential becomes 3.108 millihartree, which is a reduction of the CR-CC(2,3) NPE value by a factor of 4.4. The $N_{\text{det(in)}}=250,000$ CC($P$;$Q$) computations, which extract the lists of triples from the relatively inexpensive Hamiltonian diagonalizations in spaces that are 32–37 times smaller than the number of $T_{3}$ amplitudes used by CCSDT, reduce the 13.788 millihartree NPE characterizing the lowest-energy ${}^{1}A_{g}(D_{2\text{h}})$ potential obtained with CR-CC(2,3), relative to its CCSDT counterpart, to an impressively small value of 0.187 millihartree. This is a 74-fold reduction in NPE compared to CR-CC(2,3).
It is clear from Table 1 and Figure 2 (b) that the convergence of the lowest-energy ${}^{1}A_{g}(D_{2\text{h}})$ potentials resulting from the CIPSI-driven CC($P$;$Q$) calculations toward their CCSDT parent, including the challenging barrier region, with the CIPSI wave function termination parameter $N_{\text{det(in)}}$, with the number of determinants in the final Hamiltonian diagonalization space used to determine $|\Psi^{(\text{CIPSI})}\rangle$ [$N_{\text{det(out)}}$], and with the fraction of triply excited determinants in the $P$ space captured by CIPSI is very fast, but one cannot say the same about the uncorrected CC($P$) energies. As shown in Table 1 and Figure 2 (a), and in line with the formal analysis in Section II, the CC($P$) energies improve the CCSD results and converge toward CCSDT, but they do it at a much slower rate than their $\delta(\mbox{$P$;$Q$})$-corrected CC($P$;$Q$) counterparts. For instance, the CIPSI-driven CC($P$) computations for the lowest singlet state of cyclobutadiene using $N_{\text{det(in)}}=100,000$ reduce the 26.827, 27.964, 29.667, 32.473, 37.662, and 47.979 millihartree errors relative to CCSDT obtained at $\lambda=0$, 0.2, 0.4, 0.6, 0.8, and 1 with CCSD and the associated NPE value of 21.152 millihartree to 22.183, 23.029, 23.947, 30.198, 21.842, 32.125, and 10.283 millihartree, respectively. This should be compared to the much smaller error and NPE values characterizing the corresponding CC($P$;$Q$) calculations, which are 0.431–3.539 and 3.108 millihartree, respectively. The analogous CC($P$) calculations using $N_{\text{det(in)}}=250,000$, where the errors characterizing the $\delta(\mbox{$P$;$Q$})$-corrected CC($P$;$Q$) energies in the entire $\lambda=0$–1 region and the overall NPE relative to CCSDT are already at the level of 0.2–0.5 millihartree, produce the 17.137–25.279 millihartree errors and the NPE of 8.142 millihartree. Even with the largest $N_{\text{det(in)}}$ value considered in this study, of $10,000,000$, the 7.252–10.238 millihartree differences between the CC($P$) and CCSDT energies of the lowest ${}^{1}A_{g}(D_{2\text{h}})$ state of cyclobutadiene in the $\lambda=0$–1 region and the NPE of 2.986 millihartree that characterizes the resulting CC($P$) potential relative to its CCSDT parent remain. All of this implies that while relaxing $T_{1}$ and $T_{2}$ amplitudes in the presence of the leading triples is important, correcting the CC($P$) energies for the remaining $T_{3}$ effects, which the CC($P$) computations using the $P$ spaces generated with the help of CIPSI do not describe, is critical to reach submillihartree accuracy levels relative to CCSDT with small fractions of triples in these spaces. We observed a similar behavior in the semi-stochastic, CIQMC- and CCMC-driven,Deustua et al. (2017); Yuwono et al. (2020); Deustua et al. (2021); Chakraborty et al. (2022) and adaptiveGururangan and Piecuch (2023) CC($P$) and CC($P$;$Q$) calculations, although based on the numerical evidence that we have generated to date, the CIPSI-driven CC($P$;$Q$) methodology investigated in this study and its recently formulated adaptive analog seem to be more effective in converging the target CC (in most of our work to date, CCSDT) energetics than their semi-stochastic counterparts. This suggests that the sequences of Hamiltonian diagonalizations utilized in the CIPSI-driven CC($P$) and CC($P$;$Q$) computations and the moment expansions defining the $\delta(\mbox{$P$;$Q$})$ corrections that are used to construct excitation spaces in the adaptive CC($P$) and CC($P$;$Q$) runs are more efficient in identifying the leading higher–than–doubly excited determinants for inclusion of the underlying $P$ spaces than the CIQMC/CCMC wave function propagations, although this topic needs to be explored further and we will return to it in the future.
As shown in Table 2 and Figure 2 (c) and (d), many of the above observations apply to the CIPSI-driven CC($P$) and CC($P$;$Q$) calculations for the lowest-energy ${}^{3}B_{1g}(D_{\rm 2h})$ potential, but, given the fact that the CR-CC(2,3) approach is already very accurate in this case, producing errors relative to CCSDT that in absolute value do not exceed 60 microhartree, the CC($P$;$Q$) computations using $N_{\text{det(in)}}>1$ offer no obvious advantages over CR-CC(2,3). Nonetheless, it is reassuring that, in analogy to the $A_{g}(D_{\rm 2h})$-symmetric singlet ground state, the differences between the energies of the ${}^{3}B_{1g}(D_{\rm 2h})$ state obtained in the CC($P$) computations using CIPSI Hamiltonian diagonalizations to create lists of the leading triply excited determinants for inclusion in the underlying $P$ spaces and their CCSDT counterparts decrease as $N_{\text{det(in)}}$ increases, independent of $\lambda$. It is also encouraging that the $\delta(\mbox{$P$;$Q$})$ corrections are as effective in bringing the CC($P$) energies to a virtually perfect agreement with the parent CCSDT data as in the case of the CIPSI-driven CC($P$;$Q$) calculations for the lowest ${}^{1}A_{g}(D_{2\text{h}})$ potential using $N_{\text{det(in)}}\geq 250,000$. By inspecting the CC($P$;$Q$) column in Table 2, one may get the impression that the incorporation of the triply excited determinants identified by the CIPSI runs using increasingly large $N_{\text{det(in)}}$ values in the preceding CC($P$) steps worsens the CR-CC(2,3) results for the lowest ${}^{3}B_{1g}(D_{\rm 2h})$ state, which correspond to $N_{\text{det(in)}}=1$, but reading Table 2 in this way would be misleading. Indeed, in single-reference situations, such as that created by the lowest ${}^{3}B_{1g}(D_{\rm 2h})$ state of cyclobutadiene, where many-electron correlation effects are essentially only dynamical, the triples correction of CR-CC(2,3) [similarly to CCSD(T)] often overshoots the parent CCSDT energies (slightly). Once one starts adding triply excited determinants to the $P$ space, the CC($P$;$Q$) energies initially go up, becoming upper bounds to their CCSDT counterparts, but when the fraction of triples in the $P$ space is large enough, the differences between the CC($P$;$Q$) and CCSDT energies decrease, steadily approaching 0. We see some of this behavior in Table 2, but we have to keep in mind that the $P$ spaces used in our CIPSI-driven CC($P$;$Q$) calculations use small fractions of triples, so that the residual, $\sim$0.1 millihartree, errors relative to CCSDT remain. What is most important here is that the CC($P$;$Q$) computations for the lowest triplet potential of cyclobutadiene using $N_{\text{det(in)}}>1$ reported in Table 2 and Figure 2 (d) do not substantially alter the already excellent CR-CC(2,3) energetics. This allows us to conclude that the coupling of the lower-rank $T_{1}$ and $T_{2}$ clusters with their higher-rank $T_{3}$ counterpart is negligible in this case and the relaxation of the CCSD values of $T_{1}$ and $T_{2}$ amplitudes by including some triples in the iterative CC($P$) steps is not necessary for obtaining high-accuracy CC($P$;$Q$) results.
Having demonstrated the excellent performance of the CIPSI-driven CC($P$;$Q$) approach in accurately approximating the lowest-energy ${}^{1}A_{g}(D_{\rm 2h})$ and ${}^{3}B_{1g}(D_{\rm 2h})$ potentials of cyclobutadiene along its automerization coordinate obtained with CCSDT, in the final part of our discussion, we turn to the CC($P$) and CC($P$;$Q$) singlet–triplet gaps and their dependence on $\lambda$ and $N_{\text{det(in)}}$ examined in Table 3 and Figure 2 (e) [the uncorrected CC($P$) energetics] and (f) [the CC($P$;$Q$) results]. As pointed out above, to obtain accurate $\Delta E_{\text{S--T}}$ values for cyclobutadiene in the vicinity of the barrier region, one has to balance significant nondynamical correlations associated with the multiconfigurational singlet state, which manifest themselves in massive $T_{3}$ clusters that are strongly coupled to the one- and two-body components of $T$, with predominantly dynamical correlations characterizing the lowest triplet state that result in the generally smaller $T_{3}$ contributions having minimal effect on $T_{1}$ and $T_{2}$ amplitudes. The singlet–triplet gap values reported in Table 3 and Figure 2 (e) and (f) clearly show that neither the CCSD approach nor the CR-CC(2,3) triples correction to CCSD, which are equivalent to the CIPSI-driven CC($P$) and CC($P$;$Q$) calculations using $N_{\text{det(in)}}=1$, can do this. Both of these methods struggle with achieving a balanced description of the ${}^{1}A_{g}(D_{\rm 2h})$ and ${}^{3}B_{1g}(D_{\rm 2h})$ states of cyclobutadiene as $\lambda$ approaches 1, producing errors relative to CCSDT that are as large as 5.261 and 2.282 kcal/mol, respectively, at $\lambda=0.6$, where the CCSDT value of $\Delta E_{\text{S--T}}$ is $-10.295$ kcal/mol, 8.589 and 4.434 kcal/mol, respectively, at $\lambda=0.8$, where $\Delta E_{\text{S--T}}$ obtained with CCSDT is $-5.804$ kcal/mol, and 15.120 and 9.222 kcal/mol at $\lambda=1$, where the CCSDT result for $\Delta E_{\text{S--T}}$ is $-4.783$ kcal/mol. These large error values in the singlet–triplet gaps resulting from the CCSD and CR-CC(2,3) computations in the barrier region are a consequence of the dramatic increase in the magnitude of $T_{3}$ effects characterizing the ${}^{1}A_{g}(D_{\rm 2h})$ state and a rapidly deteriorating description of this state by both CCSD and CR-CC(2,3) as $\lambda\rightarrow 1$, seen in Table 1 and Figure 2 (a) and (b), as opposed to the nearly constant $T_{3}$ contributions and a virtually perfect agreement between the CR-CC(2,3) and CCSDT ${}^{3}B_{1g}(D_{\rm 2h})$ potentials in the entire $\lambda=0$–1 region shown in Table 2 and Figure 2 (d) [while quantitatively inaccurate, the shape of the ${}^{3}B_{1g}(D_{\rm 2h})$ potential obtained with CCSD, shown in Figure 2 (c), is qualitatively correct too]. In analogy to the previously discussed calculations for the lowest-energy ${}^{1}A_{g}(D_{\rm 2h})$ potential, in order to bring the above errors down, the input parameter $N_{\text{det(in)}}$, which controls the CIPSI diagonalization sequences preceding the CC($P$) and CC($P$;$Q$) computations, must be increased, so that the $P$ spaces used in these computations are augmented with the leading triply excited determinants, the CCSD values of $T_{1}$ and $T_{2}$ amplitudes, used in CR-CC(2,3), are properly relaxed, and the quality of the $\delta(\mbox{$P$;$Q$})$ corrections that capture the remaining $T_{3}$ effects improves.
This is precisely what we observe in the CIPSI-driven CC($P$) and CC($P$;$Q$) calculations of the singlet–triplet gaps, especially the latter ones, reported in Table 3 and Figure 2 (e) and (f). Indeed, the CC($P$;$Q$) approach using $N_{\text{det(in)}}=100,000$, which relies on small CIPSI diagonalization spaces [small $N_{\text{det(out)}}$ values], whose dimensionalities are about 1% of the $14,483,876$ $A_{g}(D_{\rm 2h})$-symmetric $S_{z}=0$ and $14,339,992$ $B_{1g}(D_{\rm 2h})$-symmetric $S_{z}=1$ triply excited amplitudes involved in the parent full CCSDT computations, and which employs even smaller $P$ spaces having only 0.1–0.6% of all triples, reduces the 0.553, 0.813, 1.300, 2.282, 4.434, and 9.222 kcal/mol differences between the CR-CC(2,3) and CCSDT $\Delta E_{\text{S--T}}$ values at $\lambda=0$, 0.2, 0.4, 0.6, 0.8, and 1 to 0.253, 0.301, 0.485, 0.784, 0.398, and 2.227 kcal/mol, respectively. When $N_{\text{det(in)}}$ is increased to $250,000$, where the numbers of determinants included in the final CIPSI diagonalizations preceding the CC($P$) and CC($P$;$Q$) steps are still only $\sim$2–3of all $T_{3}$ amplitudes used in the CCSDT calculations for the ${}^{1}A_{g}(D_{\rm 2h})$ and ${}^{3}B_{1g}(D_{\rm 2h})$ states and where the resulting $P$ spaces contain only 0.5–0.9% of all triples, the errors in the CC($P$;$Q$) $\Delta E_{\text{S--T}}$ gaps relative to their CCSDT counterparts at the above values of $\lambda$ decrease even more, to 0.112, 0.108, 0.128, 0.128, 0.152, and 0.263 kcal/mol, respectively, bringing the CC($P$;$Q$) and CCSDT results to a virtually perfect agreement, while reducing a computational effort compared to the CCSDT runs by orders of magnitude. As in the previously discussed results for the lowest singlet and triplet states of cyclobutadiene, especially for the challenging, $A_{g}(D_{\rm 2h})$-symmetric, singlet ground state, the $\delta(\mbox{$P$;$Q$})$ corrections play a major role in the observed error reductions, substantially improving the CC($P$) $\Delta E_{\text{S--T}}$ values, but, unlike in the CC($P$) calculations of the ${}^{1}A_{g}(D_{\rm 2h})$ and ${}^{3}B_{1g}(D_{\rm 2h})$ potentials, the singlet–triplet gaps obtained with the uncorrected CC($P$) approach using relatively small CIPSI diagonalization spaces and tiny fractions of all triples in the associated $P$ spaces can be quite accurate in their own right, reproducing the CCSDT values of $\Delta E_{\text{S--T}}$ across the entire $\lambda=0$–1 region to within $\sim$1–2 kcal/mol when $N_{\text{det(in)}}\gtrsim 250,000$. Clearly, this is a lot better than the 5.261, 8.589, and 15.120 kcal/mol errors relative to CCSDT obtained at $\lambda=0.6$, 0.8, and 1, respectively, with CCSD, demonstrating that the relaxation of $T_{1}$ and $T_{2}$ amplitudes in the presence of the leading $T_{3}$ contributions in the CIPSI-driven CC($P$) computations using sufficiently large $N_{\text{det(in)}}$ values results in a more balanced description of the many-electron correlation effects in the ${}^{1}A_{g}(D_{\rm 2h})$ and ${}^{3}B_{1g}(D_{\rm 2h})$ states, especially when $\lambda$ approaches 1, although, by inspecting Tables 13, we can also see that much of the improvement in the CCSD $\Delta E_{\text{S--T}}$ data offered by the CC($P$) approach originates from error cancellations between the ${}^{1}A_{g}(D_{\rm 2h})$ and ${}^{3}B_{1g}(D_{\rm 2h})$ CC($P$) energies. In the case of the CIPSI-driven CC($P$;$Q$) computations, we do not have to count on error cancellations, since both the total electronic energies of the lowest singlet and triplet states of cyclobutadiene at various values of $\lambda$ and the gaps between them rapidly converge toward their CCSDT parents with $N_{\text{det(in)}}$. The $\delta(\mbox{$P$;$Q$})$ corrections, in addition to being highly effective in improving the CC($P$) energies of the ${}^{1}A_{g}(D_{\rm 2h})$ and ${}^{3}B_{1g}(D_{\rm 2h})$ states and the associated $\Delta E_{\text{S--T}}$ values, are also very helpful in curing the nonsystematic error patterns in the singlet–triplet gaps observed in the CC($P$) calculations in Table 3 as $N_{\text{det(in)}}$ increases, where the differences between the CC($P$) and CCSDT values of $\Delta E_{\text{S--T}}$ go up and down or oscillate. This does not happen in the CC($P$;$Q$) calculations, where the resulting singlet–triplet gaps approach their CCSDT parents systematically and very fast as $N_{\text{det(in)}}$ is made larger, independent of $\lambda$. This is yet another demonstration of the ability of the CIPSI-driven CC($P$;$Q$) approach to provide a highly accurate and well-balanced description of the many-electron correlation effects characterizing the lowest-energy singlet and triplet states of cyclobutadiene along its automerization coordinate, which the conventional CC methods, such as CCSD, CR-CC(2,3), and other noniterative triples corrections to CCSD, cannot provide as one approaches the barrier region.
We conclude this section by noticing that the results reported in Tables 13 also show that the convergence of the CIPSI-driven CC($P$) and CC($P$;$Q$) energies of the lowest ${}^{1}A_{g}(D_{\rm 2h})$ and ${}^{3}B_{1g}(D_{\rm 2h})$ states of cyclobutadiene and the gap between them toward their respective CCSDT parents with $N_{\text{det(in)}}$ or $N_{\text{det(out)}}$ is faster than that characterizing the associated variational and perturbatively corrected CIPSI energetics toward the extrapolated $E_{\text{var}}+\Delta E_{\text{r}}^{(2)}$ values. This observation is consistent with our initial study announcing the CIPSI-based CC($P$) and CC($P$;$Q$) methodologies Gururangan et al. (2021) and the fact that the CC($P$;$Q$) calculations are capable of accurately approximating the parent CCSDT energetics out of the unconverged CIPSI runs using relatively small Hamiltonian diagonalization spaces, even when $T_{3}$ correlations are large and nonperturbative and electronic quasi-degeneracies become substantial, as is the case in the barrier region of the ground-state ${}^{1}A_{g}(D_{2\text{h}})$ potential. While this may not be a general remark, we also observe that perturbatively corrected $E_{\text{var}}+\Delta E^{(2)}$ and $E_{\text{var}}+\Delta E^{(2)}_{r}$ energies of the lowest singlet and triplet states of cyclobutadiene and the gap between them converge toward their extrapolated limits at a rate similar to that characterizing our uncorrected CC($P$) calculations toward CCSDT. The $\delta(\mbox{$P$;$Q$})$-corrected CC($P$;$Q$) energetics converge to CCSDT much faster. This might be yet another way of looking at the effectiveness of the $\delta(\mbox{$P$;$Q$})$ moment corrections in improving the underlying CC($P$) results. That being said, we should keep in mind that the algorithms used to obtain the CCSDT and the perturbatively corrected and extrapolated CIPSI energies are fundamentally different procedures. Furthermore, and more importantly given the objectives of this study, where we are interested in exploring the CIPSI-driven CC($P$) and CC($P$;$Q$) methodologies, not the CIPSI approach itself, the CIPSI wave function growth in the calculations reported in Tables 13 was terminated long before our CIPSI runs were well converged, as we only needed information about the leading triply excited determinants and were not interested in saturating the triply excited manifolds of the relevant many-electron Hilbert spaces. Last but not least, to highlight the robustness of our CC($P$;$Q$) framework, all of the calculations reported in this work relied on the RHF and ROHF orbitals, i.e., we made no attempt to further optimize orbitals to make them consistent with correlated computations, which would improve CIPSI’s performance and which might also help our CC($P$) and CC($P$;$Q$) results using smaller $N_{\text{det(in)}}$ values, especially in the vicinity of the square TS geometry. We intend to look into potential benefits that might be offered by orbital optimizations in the CIPSI-driven CC($P$) and CC($P$;$Q$) calculations in a future study.
IV Summary and Concluding Remarks
An accurate determination of singlet–triplet gaps in biradicals represents a formidable test for ab initio electronic structure methodologies, as it requires balancing strong nondynamical many-electron correlation effects, needed for a reliable description of the low-spin singlet states that have a manifestly multiconfigurational nature, with the generally weaker, largely dynamical, correlations characterizing the high-spin triplet states. Although high-level CC methods with a full treatment of higher–than–two-body clusters, such as CCSDT or CCSDTQ, are often powerful enough to capture the dynamical and nondynamical correlation effects relevant in such studies, their applications are hindered by the demanding computational steps and memory requirements, which are prohibitively expensive when larger many-electron systems are examined. One of the promising ideas aimed at addressing this situation within the single-reference CC framework is the CC($P$;$Q$) formalism, in which one solves the CC amplitude equations in a suitably defined subspace of the many-electron Hilbert space, referred to as the $P$ space, and improves the resulting CC($P$) energies using the a posteriori moment corrections, designated as $\delta(P;Q)$, calculated with the help of the complementary $Q$ space.Shen and Piecuch (2012, 2012, 2012); Bauman et al. (2017); Magoulas et al. (2018); Yuwono et al. (2019); Deustua et al. (2017); Yuwono et al. (2020); Deustua et al. (2021); Chakraborty et al. (2022); Gururangan et al. (2021); Gururangan and Piecuch (2023) Among the most attractive features of the CC($P$;$Q$) methodology is its flexibility, so that in addition to conventional choices of the $P$ and $Q$ spaces using truncations based on excitation ranks, which in the past resulted in the development of the biorthogonal CR-CC methods, such as the CR-CC(2,3) triples correction to CCSD, Piecuch and Włoch (2005); Piecuch et al. (2006); Włoch et al. (2006, 2007) one can consider various unconventional ways of setting up these spaces that can improve the CR-CC(2,3), CCSD(T), and similar energetics for systems with substantial electronic quasi-degeneracies by relaxing the $T_{1}$ and $T_{2}$ components of the cluster operator $T$ in the presence of their higher-rank $T_{n}$ counterparts with $n>2$, such as $T_{3}$, which become large, nonperturbative, and strongly coupled to $T_{1}$ and $T_{2}$ in such situations. This can be done without major increases in the computational effort by incorporating the leading higher–than–doubly excited determinants in the $P$ spaces and using corrections $\delta(P;Q)$ to capture the remaining correlations of interest.
In this work, we have examined the hybrid variant of the CC($P$;$Q$) methodology introduced in Ref. Gururangan et al., 2021, in which the leading higher–than–doubly excited determinants in the $P$ space are identified, in an automated fashion, using the sequences of Hamiltonian diagonalizations generated with the CIPSI algorithm.Huron et al. (1973); Garniron et al. (2017, 2019) In order to thoroughly test the CIPSI-driven CC($P$;$Q$) formalism and obtain useful insights into its performance, we have focused on recovering the lowest-energy singlet and triplet potentials of cyclobutadiene along its automerization coordinate and the gap between them resulting from the full CCSDT computations which, based on comparisons with the perturbatively corrected and extrapolated CIPSI and DEA-EOMCC(4p-2h)-level data, obtained in this study as well, provide reliable information. To do so, we have constructed an approximate, $D_{\rm 2h}$-symmetric, one-dimensional automerization pathway connecting the rectangular reactant and product species via the square TS structure on the ground-state singlet potential using the information taken from Ref. Eckert-Maksić et al., 2006 and performed a large number of CC($P$;$Q$) calculations for the lowest ${}^{1}A_{g}(D_{\rm 2h})$ and ${}^{3}B_{1g}(D_{\rm 2h})$ states of cyclobutadiene at the selected nuclear geometries along the resulting path, where for each state and for each geometry, we have explored a wide range of values of the CIPSI wave function termination parameter $N_{\text{det(in)}}$ that controls the Hamiltonian diagonalization sequences preceding the CC($P$) and CC($P$;$Q$) runs. We have demonstrated that the CIPSI-driven CC($P$;$Q$) calculations are capable of accurately approximating the high-level CCSDT energetics of the lowest singlet and triplet states of cyclobutadiene across the entire automerization pathway, to within small fractions of a millihartree for total energies and 0.1–0.2 kcal/mol for the singlet–triplet gaps, using tiny fractions of the triply excited determinants, on the order of 1% of all triples, in the underlying $P$ spaces extracted from the relatively inexpensive CIPSI diagonalizations in spaces that are orders of magnitude smaller than the numbers of cluster amplitudes used by CCSDT. This extraordinary performance of the CIPSI-driven CC($P$;$Q$) approach applies to both the less demanding reactant/product region, where the lowest singlet and triplet states of cyclobutadiene are largely single-configurational and the absolute values of the singlet–triplet gap exceed 30 kcal/mol, and the vicinity of the TS structure on the ground-state singlet potential, where the high-spin triplet state retains its weakly correlated, single-determinantal, nature, but the singlet state, separated from its triplet counterpart by only $\sim$5 kcal/mol, becomes multiconfigurational, strongly correlated, and characterized by large and highly nonperturbative $T_{3}$ correlations, which are strongly coupled to the one- and two-body components of $T$ and which result in failure of CR-CC(2,3) and other noniterative triples corrections to CCSD. Interestingly, the uncorrected CC($P$) computations using similarly compact excitation spaces can be accurate as well, reproducing the CCSDT values of the singlet–triplet gap across the entire automerization pathway to within $\sim$1–2 kcal/mol, improving the poor CCSD and CR-CC(2,3) results in the vicinity of the TS geometry, and reaffirming the usefulness of incorporating the leading triply excited determinants into the underlying $P$ spaces, but the total energies of the lowest ${}^{1}A_{g}(D_{\rm 2h})$ and ${}^{3}B_{1g}(D_{\rm 2h})$ states of cyclobutadiene obtained with the CC($P$) approach converge to their CCSDT parents with $N_{\text{det(in)}}$ very slowly, so one has to rely on error cancellations to obtain accurate singlet–triplet gaps with CC($P$). The $\delta(\mbox{$P$;$Q$})$ corrections are very helpful in this regard. They reduce errors in the total CC($P$) energies of the ${}^{1}A_{g}(D_{\rm 2h})$ and ${}^{3}B_{1g}(D_{\rm 2h})$ states of cyclobutadiene by orders of magnitude while substantially improving the resulting singlet–triplet gaps, making their convergence toward CCSDT smoother and more systematic. Given the relatively low costs of determining the $\delta(\mbox{$P$;$Q$})$ corrections compared to the preceding CIPSI and CC($P$) steps and the enormous benefits resulting from their application in the CIPSI-driven CC($P$;$Q$) calculations, we recommend using CC($P$;$Q$).
The excellent performance of the CIPSI-driven CC($P$;$Q$) approach in converging the lowest-energy singlet and triplet potentials of cyclobutadiene obtained with CCSDT, observed in this study, along with the promising initial results reported in Ref. Gururangan et al., 2021, motivate us to pursue the hybrid CC methodologies combining the CC($P$;$Q$) framework with selected CI even further. We will, for example, investigate how much the CC($P$;$Q$) singlet and triplet potentials reported in this work, especially the ground-state singlet potential in the vicinity of the TS geometry, can benefit from replacing the RHF and ROHF orbitals exploited in the calculations reported in this work by the suitably optimized orbitals consistent with the CC($P$) or CIPSI wave functions. We will also examine how much each of the singlet and triplet potentials of cyclobutadiene obtained with the CIPSI-driven CC($P$;$Q$) approach using a given value of $N_{\text{det(in)}}$, especially its smoothness, can improve by consolidating the $P$ spaces corresponding to the different geometries along the automerization path or, to be more precise in the context of the CC($P$;$Q$) calculations aimed at recovering the CCSDT energetics performed in this study, by merging the triple excitation manifolds incorporated in those spaces. Among other topics worth exploring, it will be interesting to examine if anything substantial can be gained by replacing the CIPSI algorithm in our CC($P$;$Q$) considerations by other selected CI techniques, such as heat-bath CI,Holmes et al. (2016); Sharma et al. (2017); Li et al. (2018) adaptive CI,Schriber and Evangelista (2016, 2017) or adaptive sampling CI.Tubman et al. (2016, 2020) Last but not least, following the strategy adopted in our previous work on the semi-stochastic, CIQMC-driven, CC($P$;$Q$) approaches, Deustua et al. (2017); Yuwono et al. (2020); Deustua et al. (2021) we are planning to extend the CIPSI-driven CC($P$;$Q$) methodology investigated in this study and Ref. Gururangan et al., 2021, to higher CC levels, especially CCSDTQ, and excited electronic states, with an initial focus on converging the EOMCCSDTKowalski and Piecuch (2001, 2001); Kucharski et al. (2001) energetics, while seeking additional savings in the computational effort by replacing the unconstrained CIPSI algorithm, which is allowed to explore the entire many-electron Hilbert space, by its truncated analogs consistent with the determinantal spaces needed in the target CC calculations (e.g., the CISDT or CISDTQ analogs of CIPSI when attempting to use the CIPSI-driven CC($P$;$Q$) framework to converge the CCSDT or CCSDTQ energetics).
Acknowledgements.
This work has been supported by the Chemical Sciences, Geosciences and Biosciences Division, Office of Basic Energy Sciences, Office of Science, U.S. Department of Energy (Grant No. DE-FG02-01ER15228 to P.P.).
Notes
The authors declare no competing financial interest.